OpenStudy (anonymous):

A wire runs from the top of a flagpole to a point on the ground that is 9 yards from the base of the pole. The wire is 15 yards long. How many yards tall is the flagpole?

OpenStudy (ganpat):

can u find x , by Pythagoras theorem ?? 

OpenStudy (ganpat):

hellloooo ??

OpenStudy (ganpat):

Pythagoras ?? heard anywhere ??

OpenStudy (ganpat):

x2 + 9square = 15 square x2 = 225 -81 x2 = 144.. so x = 12... i.e height of the pole...